### v2

# Searching Meta Reasoning Skeleton to Guide LLM Reasoning

Abstract
Meta reasoning behaviors work as a skeleton to guide large language model (LLM) reasoning, thus help to improve reasoning performance. However, prior researches implement meta reasoning skeleton with manually designed structure, limiting ability to adapt to query-specific requirement and capture intricate logical dependency among reasoning steps. To deal with the challenges, we represent meta reasoning skeleton with directed acyclic graph (DAG) to unify skeletons proposed in prior works and model intricate logical dependency. Then we propose AutoMR, a framework that searches for query-aware meta reasoning skeleton automatically inspired by automated machine learning (AutoML). Specifically, we construct search space based on DAG representation of skeleton and then formulate the search problem. We design a dynamic skeleton sampling algorithm by expanding meta reasoning skeleton along with reasoning context at inference time. This algorithm can derive any meta reasoning skeleton in search space efficiently and adapt skeleton to evolving base reasoning context, thus enable efficient query-aware skeleton search. We conduct experiments on extensive benchmark datasets. Experimental results show that AutoMR achieves better reasoning performance than previous works broadly.
1 Introduction
Large language model (LLM) demonstrate superior performance on complex tasks such as math Q&A when equipped with step-by-step reasoning ability (cot; o1; r1). Researches on cognition divide reasoning into two levels: base reasoning (reasoning for problem directly) and meta reasoning (higher-level reasoning about how to reason) (metacognition). Meta reasoning, considered a unique ability of human cognition (ackerman2017meta), entails awareness of one’s reasoning process and the deliberate selection of reasoning strategies. For instance, when encountering difficulty with math problem, humans shift solution by thinking “This approach is not working; I should try another method…” or they may verify their reasoning steps by reflecting “Some steps may have errors. Let me check a previous step…” These behaviors do not directly solve the problem itself but instead organized as skeleton to guide the reasoning process.
Inspired by such human behaviors, previous studies proposed to incorporate meta reasoning into LLM to guide their reasoning process and thereby enhance performance on complex reasoning tasks (mrp; rstar; metareasoner; metascale). Recent approaches typically predefine a set of meta reasoning strategies for intermediate reasoning steps and employ manually designed structures (e.g. sequential, parallel and tree) to organize the strategies into meta reasoning skeleton. For example, rStar (rstar) and Meta-Reasoner (metareasoner) both define step-wise strategies such as decomposing question into sub-questions.
Figure 1: Human behaviors in meta reasoning for three questions about math (Q1 and Q2) and biology multi-choice (Q3).
rStar leverages Monte Carlo Tree Search (MCTS) (mcts) to select and organize strategies, whereas Meta-Reasoner arranges them in a sequential way and selects at each step via multi-armed bandit (mab). An intuitive illustration of these manually designed skeleton is provided in Figure 2.
The aforementioned methods based on manually designed meta reasoning skeleton improved LLM reasoning performance. However, evidence from cognition science suggests that meta reasoning skeletons should vary for different queries, due to reasoner ability, query difficulty, discipline characteristic, etc. (scott2013examining; erickson2015metacognition; rouault2018human). For example in Figure 1, knowledge-intensive problems (Q3 about biology) rely more heavily on knowledge-recall strategy while shallower thinking depth than thinking-intensive problem (Q1 and Q2 about math). More difficult problems (Q1) may demand more parallel reasoning branches with solution exploration strategy than easier one (Q2). Besides, the logical dependency of reasoning steps can be too intricate (besta2024graph) to capture by sequential, parallel, or tree-structured skeletons in prior works. The skeleton of Q1 involves parallel branches (steps 1–3 forming one branch while steps 4–6 another) and multiple dependency (step 6 simultaneously depends on step 5 as well as step 3 from early branches). Skeleton of Q3 summarizes two steps (step 2 and 3) to make it answer confident. The query-specific requirement and the intricate logical dependency among reasoning steps make it challenging for existing methods with limited manually designed meta reasoning skeletons (Figure 2) to work well across all queries.
Automated machine learning (AutoML) seeks to generate machine learning configurations for given task in a data-driven manner (automl), thereby reducing the need for manual design and tuning for neural architectures (nas) and hyperparameter (feurer2019hyperparameter). Inspired by success of AutoML, we propose AutoMR, a framework that automatically searches for query-aware meta reasoning skeletons to guide LLM to reason for correct answer, where we represent meta reasoning skeleton as single-source edge-heterogeneous directed acyclic graph (DAG) to cover skeleton in prior works and capture intricate logical dependencies. Specifically, we first design an extensive DAG-based skeleton search space. Then we formulate the meta reasoning skeleton search problem, which poses two technical difficulties specific to query-aware skeleton search. The first is to derive any skeleton for given query from the extensive search space efficiently. The other is to adapt derived skeleton to evolving base reasoning context, considering inherent step-by-step property of reasoning process. To tackle the difficulties, we design a skeleton sampling algorithm that expands meta reasoning skeleton node by node dynamically based on base reasoning context at inference time. We prove that this algorithm introduces minimal additional computation overhead compared with naive LLM reasoning process. Compared with prior meta reasoning method, our search for meta reasoning skeleton improves reasoning performance. Moreover, we show that our search and inference algorithm is efficient theoretically and empirically.
We summarize our contributions as follows:
  • We propose AutoMR to search for query-aware meta reasoning skeleton, where we represent meta reasoning skeleton as DAG to capture intricate logical dependency among reasoning steps.
  • We design an extensive skeleton search space based on DAG. Additionally, we introduce an dynamic skeleton sampling algorithm that can derive any skeleton in search space efficiently and adapt skeleton to evolving base reasoning context at inference time.
  • We conduct experiments on benchmark datasets across different disciplines and difficulties. Experimental results show that AutoMR demonstrates better reasoning performance than previous meta reasoning methods, with high search and inference efficiency.
2 Related Works
Meta Reasoning in LLM. Meta reasoning is an ability of human cognition involving determining reasoning strategy about how to reason (metacognition; ackerman2017meta). Previous works explored to introduce meta reasoning into LLM to guide it reasoning (metascale; alazraki-rei-2025-meta; yan2025position; metacot; de2024rational; rema; didolkar2024metacognitive). Meta Reasoning Prompt (MRP) (mrp) includes classic strategies like CoT (cot), Self-Refine (selfrefine), etc. It first prompts LLM to choose one strategy for given query and then reason guided by that strategy. Strategies in MRP are holistic, meaning that MRP uses only one strategy for the whole reasoning process without adjusting when reasoning progressing. In contrast, recent methods usually use step-wise meta reasoning strategies (reasonflux; yang2025supercorrect) and choose strategy for each step during reasoning. For example, rStar (rstar) define step-wise reasoning strategies such as proposing a sub-question, and then use MCTS to build tree-structured meta reasoning skeleton. Meta-Reasoner (metareasoner) also uses step-wise reasoning strategies but organizes them with sequential skeleton and uses multi-armed bandit to select strategy for each step. This kind of methods incorporate more fine-grained meta reasoning guidance and allow adjusting strategies during reasoning, thus performing better empirically than MRP.
Automated Machine Learning (AutoML). AutoML aims to search for high-performing machine learning (ML) configuration for given task automatically, reducing demand for human manual design (he2021automl) to adapt to task-specific requirement. Typical AutoML atomizes ML configurations to construct search space and develop search algorithm to find effective candidates (automl). Previous works implemented this idea for multiple ML configurations such as neural architecture search (NAS) (nas1000; darts; enas) and hyperparameter search (yang2020hyperparameter; shen2023efficient), and have achieved success. For example, architectures found by NAS surpass human-designed ones on various tasks, such as computer vision (real2019regularized) and natural language processing (so2019evolved). Recent works explored integrating AutoML with LLMs, like automating LLM agent workflow building (zhuge2024gptswarm; zhang2025multiagent; saad-falcon2025an). However, applying AutoML method to search for meta reasoning skeleton is non-trivial due to factors specific to LLM reasoning task, including query-specific requirement, intricate logical dependency, and evolving reasoning context.
3 Proposed Method
We introduce AutoMR that automatically searches for query-aware meta-reasoning skeletons to guide LLM reasoning. Section 3.1 presents a unified perspective on meta-reasoning skeleton in existing meta-reasoning methods based on DAG to capture intricate logical dependency. With this unified view, we construct our skeleton search space. Section 3.2 formulates the meta-reasoning skeleton search problem and details our overall search strategy. Finally, Section 3.3 discusses comparison with techniques in AutoML and analyzes our advantage specific to LLM reasoning tasks.
3.1 Search Space
Given a query $q$, let $\mathcal{S}$} denote the set of meta reasoning strategies for intermediate reasoning steps. The objective of a meta-reasoning method is to organize strategies from $\mathcal{S}$} into meta reasoning skeleton to direct LLM on performing reasoning to answer $q$.
Prior works use manually designed meta reasoning skeleton structure (e.g. sequential, parallel, tree-structured in Figure 2). To unify these designs and capture intricate logical dependencies (Figure 1), we represent meta reasoning skeleton as a single-source, edge-heterogeneous directed acyclic graph (DAG). Formally, a meta reasoning skeleton can be represented as a DAG $\alpha=({\mathcal{V}},{\mathcal{E}},\tau,{\mathcal{S}})$. Node $n_{i}=(i,c_{i})\in{\mathcal{V}}$ representing a reasoning step, $i$ being the topological index and $c_{i}$ textual content of the step. Edge $(i,j)\in{\mathcal{E}}$ indicating reasoning progression from $n_{i}$ to $n_{j}$. $\tau:{\mathcal{E}}\rightarrow{\mathcal{S}}$ maps edge to its strategy, under which LLM generates the reasoning text. There exists a unique source node $n_{0}$ with $c_{0}=q$, making $\alpha$ single-source. With above representation, we have Proposition 1 to cover the skeletons in prior works. See Appendix B.1 for proof.
Proposition 1.
Sequential, parallel, and tree structured skeletons can all be represented as single-source, edge-heterogeneous DAGs.
Figure 2: Overview of the AutoMR. Top: Illustration of search space, an example skeleton sampling process and resulting sampled skeleton. Node 0 is the single source node representing query. Steps (1)(2)(3) show how nodes 1, 2, and 3 are successively added to partial skeleton. For clarity, we display only 4 nodes and 2 types of meta reasoning strategies (red and blue edges), and the zero option (gray edges); In practice, the number of nodes can be arbitrary if token budget is satisfied and we actually implement richer strategies. Bottom: Search space subsumes sequential, parallel, and tree-structured skeletons.
Based on this unified view, we construct search space to contain all skeletons represented by single-source edge-heterogeneous DAG as shown in Figure 2, as long as the sum of tokens for all node content except source node (i.e. number of tokens generate by LLM) does not reach token budget $\mathcal{B}$}, where $\mathcal{B}$} is a hyperparameter.
We summarize the meta reasoning behaviors in previous works about LLM reasoning (gandhi2025cognitivebehaviorsenableselfimproving; chen2025reasoningerasurveylong), which gives meta reasoning strategy set ${\mathcal{S}}=\{\textbf{Next},\textbf{Reflect},\textbf{Explore},\textbf{Decompose},\textbf{Summarize},\textbf{Recall},\textbf{Answer}\}$. All of these meta reasoning strategies are implemented by designed prompt. Functions and prompt of these strategies are summarized in Table 3 in Appendix A.1. Following previous works (darts), we also introduce a special zero edge type to indicate an edge in fact does not exists.
Given meta reasoning strategy set $\mathcal{S}$} and token budget $\mathcal{B}$}, search space $\mathcal{A}$} is defined as follows,
${\mathcal{A}}=\Big\{\alpha=({\mathcal{V}},{\mathcal{E}},\tau,{\mathcal{S}})\mid\alpha\text{ is single-source DAG},\;\;\tau:{\mathcal{E}}\rightarrow{\mathcal{S}},\;\;\sum\nolimits_{n_{i}\in{\mathcal{V}}\setminus\{n_{0}\}}|c_{i}|\leq{\mathcal{B}}\,\Big\},$ (1)
where ${\mathcal{V}}\setminus\{n_{0}\}$ is node set without $n_{0}$ and $|c_{i}|$ denote number of tokens in content $c_{i}$. As illustrated in Figure 2 (bottom), this search space includes all single-source DAGs, thus subsuming skeletons considered in prior meta-reasoning methods, such as sequential, parallel, and tree-structured forms.
3.2 Search Strategy
Next, we now provide the formal definition of meta-reasoning skeleton search problem. Considering that the meta-reasoning skeleton should depend on the specific query (e.g., query difficulties and discipline characteristics), the problem is formulated as follows.
Definition 1 (Meta-Reasoning Skeleton Search Problem).
Let $\mathcal{S}$} denote meta reasoning strategy set and $\mathcal{A}$} the skeleton search space defined on $\mathcal{S}$}. $(q,a)$ is query–answer pair from dataset $\mathcal{D}$}. Given policy $P$ that derives a meta reasoning skeleton $\alpha_{q}\in{\mathcal{A}}$ for query $q$, the search objective is
$\operatorname*{arg\,max}\nolimits_{P}\mathbb{E}_{(q,a)\sim{\mathcal{D}},\alpha_{q}\sim P(\cdot|q)}[r(a,\text{LLM}(q;\alpha_{q}))].$ (2)
Here $\text{LLM}(q;\alpha_{q})$ denotes LLM reasoning on query $q$ under guidance of $\alpha_{q}$, and $r$ measures reasoning performance against the ground-truth answer $a$.
When implementing a policy $P$ for deriving a query-aware skeleton, this search problem poses two technical challenges specific to LLM reasoning. First, the search space is extensive, so the derivation procedure must efficiently explore it to recover arbitrary skeletons in it. Second, because reasoning process unfolds step by step (cot; nye2022show), the derivation process should adapt meta reasoning strategy at each step in skeleton to evolving base reasoning context, rather than fixing the skeleton a priori before reasoning for given query.
To address above difficulties, Section 3.2.1 introduces a skeleton-sampling algorithm that expand skeleton node by node dynamically, along with base reasoning context at inference time. We prove that the algorithm can cover any skeleton in search space within minimal additional computation compared with naive LLM reasoning process; Section 3.2.2 presents the overall search algorithm.
3.2.1 Dynamic Skeleton Sampling at Inference Time
We introduce an efficient algorithm that sample skeleton dynamically to implement policy $P(\cdot\mid q)$. Considering step-by-step nature of reasoning, step-wise meta reasoning strategy should adapt to current base reasoning context. This makes it necessary to interleave meta reasoning with base reasoning. To realize this, we sample skeleton starting from the single source node as a partial skeleton, and then expand it node by node in topological order, dynamically align with step-by-step base reasoning at inference time .
Algorithm 1 Dynamic Skeleton Sampling at inference time
0: Query $q$, token budget $\mathcal{B}$}
0: Meta reasoning architecture $\alpha_{q}$
1: Initialize $\alpha_{q}$ as empty DAG, $i\leftarrow 0$
2:while $\mathcal{B}$} is not reached do
3:  for $j$ from ii-1 to 0 do
4:   Sample $s_{(j,i)}\sim p_{\theta}(s_{(j,i)}|c_{j},s_{(>j,i)},c_{:i-1})$ with MLP
5:  end for
6:  if all sampled strategies are zero then
7:   Generate final answer and return
8:  end if
9:  Generate content $c_{i}$ for $n_{i}$, $i\leftarrow i+1$
10:end while
11: Generate final answer
Specifically, we set content $c_{0}$ of $n_{0}$ as $q$, forming a partial architecture. Expansion then proceeds in topological order. For each target node $n_{i}$, we determine the existence and types of incoming edges before (optionally) generating its content. Concretely, when visiting $n_{i}$ we first activate it (no content yet) and perform following three steps.
Step1: Determine incoming edges for meta reasoning. Traverse existing nodes $n_{j}$ ($0\leq j\leq i-1$) in reverse order (from $n_{i-1}$ to $n_{0}$) and sample a strategy $s_{(j,i)}\in{\mathcal{S}}\cup\{\textit{zero}\}$ for each potential edge $(j,i)$. Each sampling is conditioned on the predecessor content $c_{j}$, the already chosen strategies $s_{(>j,i)}$ for $n_{i}$, and the current base reasoning context $c_{:i-1}$ (the contents of $n_{0},\dots,n_{i-1}$), which is computed as $p(s_{(j,i)}|c_{j},s_{(>j,i)},c_{:i-1})$.
Step2: Check completion. If all sampled strategies are zero (no edge enters $n_{i}$), we deem the skeleton complete without adding $n_{i}$ and prompt the LLM to produce the final answer from the current context $c_{:i-1}$.
Step3: Generate base reasoning content. If at least one incoming edge exists, we prompt the LLM under the guidance of the sampled strategies $s_{(<i,i)}$ (excluding zero) and the contents of $n_{i}$’s predecessors to produce the next base reasoning step; the generated text is assigned to $c_{i}$, and $n_{i}$ (with its incoming edges) is added as a node with content.
Then we repeat this expansion for $n_{i+1}$ until Step 2 triggers or token budget is reached.
We implement $p(\cdot)$ with a multi-layer perception (MLP) parameterized with $\theta$. The MLP takes representations of $c_{j}$, $s_{(>j,i)}$, and $c_{:i-1}$ as input and outputs logits followed by softmax to obtain distribution over ${\mathcal{S}}\cup\{\textit{zero}\}$. These representations are cached byproducts of the ongoing LLM inference (i.e. pooled hidden states), thus requiring no additional LLM calls. If the sampled skeleton $\alpha_{q}$ contains $|{\mathcal{V}}|$ nodes, its policy (also parameterized with $\theta$ now) log-probability factorizes as
$\log P_{\theta}(\alpha_{q}|q)=\sum\nolimits_{i=1}^{|{\mathcal{V}}|-1}\sum\nolimits_{j=0}^{i-1}\log p_{\theta}(s_{(j,i)}|c_{j},s_{(>j,i)},c_{:i-1}).$ (3)
The sampling process is shown in Figure 2 and formalized in Algorithm 1. According to Algorithm 1, meta reasoning strategy sampling is conditioned on current base reasoning context at each step, thereby yielding a query-aware architecture since reasoning context traces back to $c_{0}=q$. Implementation details are in Appendix A.2. For Algorithm 1, we have Proposition 2.
Proposition 2.
Algorithm 1 can derive any $\alpha\in{\mathcal{A}}$, within $O(|{\mathcal{V}}|^{2})$ additional MLP calls (line4) compared with naive LLM reasoning process.
The time complexity of naive LLM reasoning process is proportional to ${\mathcal{B}}^{2}$. But $|{\mathcal{V}}|\ll{\mathcal{B}}$ because one step usually contains many tokens, and MLP uses much less computation than LLM, so AutoMR introduces minimal additional computation relative to naive LLM reasoning. We provide proof of Proposition 2 and detailed efficiency analysis in Appendix B.2.
3.2.2 Overall Search Algorithm
With $P_{\theta}(\alpha_{q}|q)$ defined in (3), we follow REINFORCE (reinforce; zoph2017neural), a policy gradient algorithm implementing unbiased empirical approximation of objective, to optimize $\theta$. Specifically, we sample batches with $N$ query-answer pairs $(q_{i},a_{i})$ from training set each time and optimize $\theta$ with these batches iteratively. For each $(q_{i},a_{i})$ in batch, we sample $M$ skeletons $\alpha_{q_{i}}^{j}$ from $P_{\theta}(\cdot|q_{i})$ and evaluate their performance with $r(\cdot)$ respectively. The update to $\theta$ in each iteration by estimated policy gradient with a batch is as follows, where $\eta$ is learning rate:
$\theta\leftarrow\theta+\frac{\eta}{MN}\sum\nolimits_{i=1}^{N}\sum\nolimits_{j=1}^{M}[r(a_{i},\text{LLM}(q_{i},\alpha_{q_{i}}^{j}))\nabla_{\theta}\log P_{\theta}(\alpha_{q_{i}}^{j}|q_{i})].$ (4)
The overall search algorithm and implementation is provided in Appendix A.3. We do not tune LLM parameters directly, thus enabling efficient search. For inference, we follow Algorithm 1 for each query to sample meta reasoning skeleton, generate base reasoning and output final answer.
3.3 Technical Comparison with AutoML
Different from prior meta reasoning methods that rely on manually designed skeleton (rstar; metareasoner), AutoMR draws inspiration from AutoML to search for query-aware meta reasoning skeleton from DAG-based search space, thereby addressing query-specific requirements. Technically, AutoMR is related to topics in AutoML such as neural architecture search. Recent studies have extended AutoML ideas to LLM-related tasks, such as automating agent workflow building (zhuge2024gptswarm; zhang2025multiagent). However, the unique properties of LLM reasoning tasks make AutoMR particularly suited for meta reasoning skeleton search. First, reasoning queries often exhibit highly specific demands, making a single meta reasoning skeleton insufficient. Second, the reasoning process typically involves intricate logical dependencies. Third, reasoning unfolds step by step, with the base reasoning context dynamically evolving as each new step is generated. These characteristic fundamentally differs from those of neural architecture or agent workflow, which is usually fixed for all queries or static during inference. For example, Prior approaches (zoph2018learning; zhuge2024gptswarm) generally output a single architecture or agent workflow for all queries. While instance-aware methods (cheng2020instanas; zhang2025multiagent) produce input-specific architecture or workflow that remain static during inference. Such differences in task properties makes the search techniques in these methods perform well in their target scenarios but cannot be applied to meta reasoning skeleton search directly. We compare these search techniques empirically by ablation study in Section 4.3.
4 Experiments
4.1 Setup
Baselines. We implement the following types of baselines: (1) Classic methods, including Direct-I/O and CoT (cot). (2) Meta reasoning methods, including MRP (mrp), rStar (rstar) and Meta-Reasoner (metareasoner). We also include MaAS (zhang2025multiagent), a method using NAS technique to automate multi-agent workflow building.
AutoMR and all the baselines are implemented based on two LLMs including LLaMA3.2-3B-Inst (hereinafter referred to as ”LLaMA”) (llama3herdmodels) and Qwen2.5-3B-Inst (hereinafter referred to as ”Qwen”) (qwen25) to avoid impact on experimental results caused by unique properties of specific LLM (gandhi2025cognitivebehaviorsenableselfimproving). We set the same token budget to 1024 for all methods to ensure fair comparison. More implementation details of the baselines are introduced in Appendix C.1.
Datasets and Metric. We evaluate AutoMR and baselines on two domains, i.e. math Q&A and general multiple-choice. For math Q&A, we choose GSM8K (cobbe2021gsm8k), MATH-500 (hendrycksmath2021), AMC (including AMC 2022 and AMC 2023) and Olympiad (only open-ended text-only math subset to avoid influence from multi-modal and multilingual input information) (he2024olympiadbench) to evaluate. We use training split of MATH dataset to train AutoMR and baselines that need training. For general multiple-choice, we choose MMLU-Pro (wang2024mmlupro) and split it into four subsets as Science, Humanities, Social and Other referring to zhang2025right, to evaluate. We collect training split of MMLU-Pro to train. Details of these datasets are summarized in Appendix C.2. We use Accuracy as metric to evaluate these methods .
4.2 Performance Comparison
We report the overall performance of AutoMR and baselines on math Q&A datasets and general multiple-choice datasets (Table 1). Across both domains and model backbones, AutoMR consistently achieves the best results, highlighting its broad effectiveness. Our findings can be summarized as follows: (1). Effectiveness of meta reasoning methods. Meta reasoning approaches (MRP, Meta-Reasoner, rStar, and AutoMR) consistently outperform the standard CoT baseline. Notably, Meta-Reasoner—despite adopting the same sequential organization as CoT—achieves a substantial improvement, underscoring the benefits of incorporating meta reasoning behaviors. (2). Importance of fine-grained meta reasoning strategies. Among meta reasoning methods, those that leverage strategies for guiding intermediate reasoning steps (Meta-Reasoner, rStar, and AutoMR) outperform MRP, which relies on holistic strategy. This result highlights the advantage of fine-grained meta-level guidance during reasoning. (3). Advantage of DAG-based search space. Compared with Meta-Reasoner and rStar, which rely on manually designed sequential and tree-structured skeleton respectively, AutoMR achieves superior performance. (4). AutoMR surpasses automatic agent workflow MaAS, demonstrating that AutoMR is more proper for LLM reasoning tasks.
Table 1: The overall performance on math Q&A and general multi-choice. Letters after method names means the used skeleton structure. S: Sequential; T: Tree; G: DAG; “-” means not applicable.
Method MATH-500 GSM8K AMC Olympiad
LLaMA Qwen LLaMA Qwen LLaMA Qwen LLaMA Qwen
Direct-I/O (-) 12.6 16.8 11.1 15.8 12.0 8.4 3.7 5.5
CoT (S) 36.8 61.6 71.1 85.3 21.2 34.9 11.9 26.2
MRP (-) 40.8 63.8 74.6 88.2 25.3 33.7 11.6 26.6
Meta-Reasoner(S) 44.4 65.4 76.8 87.0 26.5 36.1 13.1 27.4
rStar (T) 46.6 67.0 78.9 88.7 15.7 32.5 15.1 25.4
MaAS (S) 46.2 63.6 76.4 86.4 24.1 33.7 12.6 27.7
AutoMR (G) 50.2 69.6 81.9 91.5 30.1 38.6 17.4 30.4
Method Science Humanities Social Other
LLaMA Qwen LLaMA Qwen LLaMA Qwen LLaMA Qwen
Direct-I/O (-) 16.3 32.7 11.5 25.1 15.8 39.0 14.5 29.1
CoT (S) 31.5 41.6 22.4 28.3 37.3 51.5 31.3 39.8
MRP (-) 36.4 42.8 24.2 30.1 40.6 53.5 32.8 41.6
Meta-Reasoner (S) 44.3 45.4 30.6 31.9 47.2 55.0 36.4 42.2
rStar (T) 42.6 43.6 30.0 30.8 46.8 55.4 34.8 36.0
MaAS (S) 44.6 45.5 29.7 31.0 46.2 56.0 35.6 41.7
AutoMR (G) 48.9 49.4 33.2 33.7 51.0 57.4 38.8 45.6
4.3 Ablation Study
Influence of token budget scaling. Previous works shows that LLM reasoning performance improves whentoken budget increases (o1; tts). We evaluate the performance when scaling token budget $\mathcal{B}$}. We compare AutoMR with baselines able to scale token budget. Specifically, for CoT we implement sequential scaling technique Budget Forcing (s1) and parallel technique Majority Voting (sc). We also choose Meta-Reasoner and rStar as baselines. We do not include MaAS as baselines to evaluate because it do not provide scaling technique in original paper. The scaling technique implementation details of these methods are in Appendix C.1. We evaluate on MATH-500 and Science based on Qwen. According to results in Figure 3, we observe that when token budget increases, each method improves performance on the whole. Specifically, the scaling efficiency on knowledge-intensive Science subset is much slower than that on thinking-intensive MATH-500, according with recent research (zhao2025test). Forcing sequential scaling (i.e. Budget Forcing and Meta-Reasoner) scale slowly. Majority Voting based on parallel skeleton and rStar based on tree-structured skeleton scale more efficiently than sequential ones. AutoMR achieve the highest scaling efficiency, because search space based on DAG in AutoMR allows more extensive skeleton exploration.
Figure 3: The scaling curve of AutoMR and baselines.
Figure 4: The training and inference cost and performance of AutoMR and baselines.
Effectiveness of search strategy. We evaluate the effectiveness of search strategy in Section 3.2 against Random Search (RS(randomsearch), a common AutoML baseline (li2020random). We also assess effectiveness of dynamic skeleton sampling algorithm by comparing it with two variants. Query-Invariant (QI), sampling single meta reasoning skeleton shared by all queries of a task, as in prior NAS methods (darts; enas). Complete in Advance (CA), sampling query-specific skeletons before reasoning starts but not based on reasoning context (cheng2020instanas; zhang2025multiagent). Implementation details of these sampling methods are in Appendix C.1. We compare them on MATH-500 and Science.
Table 2: Ablation study on search strategy.
Method
MATH-500 Science
LLaMA Qwen LLaMA Qwen
RS 36.2 59.4 38.5 43.3
QI 37.2 60.2 37.3 43.9
CA 50.0 66.2 45.7 47.1
AutoMR 50.2 69.6 48.9 49.4
According to results in Table 2, AutoMR achieves the best performance compared with three variants, showing the effectiveness of proposed search strategy. In terms of skeleton sampling algorithm, AutoMR and CA both surpass QI, showing the importance of query-specific meta reasoning skeleton. Moreover, AutoMR performs better than CA, demonstrating the effectiveness dynamic skeleton sampling algorithm based on evolving reasoning context compared with the complete skeleton in advance.
Training and inference efficiency. To support theoretical analysis in Section 3.2.2 that AutoMR incurs minimal additional computation, we evaluate both training and inference costs of AutoMR and baselines requiring training, including Meta-Reasoner and MaAS, based on both Qwen and LLaMA on MATH-500 dataset. We also implement GRPO (deepseekmath), a reinforcement learning method to enhance LLM reasoning, based on LoRA (lora) as a baseline in our experiment setting. Results in Figure 4 show training cost (x-axis), performance on MATH-500 (y-axis), and inference cost (circle area). In terms of training, AutoMR and other two baselines require far less time than GRPO, which fine-tunes LLM parameters directly. However, only AutoMR achieves comparable performance with Qwen and even surpasses it with LLaMA. In terms of inference, AutoMR is slightly slower than naive reasoning process based on GRPO-trained LLM and slightly faster than MaAS, while being substantially more efficient than Meta-Reasoner, which relies on additional LLM calls to summarize reasoning progress. Instead, AutoMR employs a lightweight MLP to process representations produced during reasoning, avoiding extra LLM calls.
4.4 Case Study
Figure 5: Searched skeletons for queries from MATH-500 Level1, Level5 and Science respectively.
We visualize searched meta reasoning skeletons of three queries respectively in Figure 5. Q1 and Q2 come from MATH-500 while Q3 is from Science. According to three skeletons and their corresponding queries, we observe that AutoMR can search out query-aware skeleton, which is appropriate for given query considering query properties such as difficulty and discipline characteristics.
Skeleton Cases of Queries from Different Tasks. Q1 and Q2 correspond to math Q&A tasks, which are typically regarded as thinking-intensive, while Q3, drawn from the Science subset, concerns the history of biology and is considered knowledge-intensive. For two math queries, skeletons sampled by AutoMR exhibit deeper reasoning steps and employ more diverse meta reasoning strategies (e.g., Exploration and Reflection) than that sampled for Q3. By contrast, skeleton for Q3 emphasizes Recall strategy. This distinction aligns with the characteristics of thinking-intensive math versus knowledge-intensive history of biology.
Skeleton Cases of Queries with Different Difficulties. Both Q1 and Q2 are drawn from the MATH-500 dataset, Q2 belongs to the more challenging “Level-5” subset whereas Q1 comes from simpler “Level-1” subset. Correspondingly, skeleton for Q1 is more complex than that of Q2. In Figure 5, the skeleton for Q1 contains two reasoning branches, where the LLM explores two potential solutions, with the first attempt failing. It also incorporates Recall strategy to leverage intermediate result from earlier steps. However, skeleton for simpler Q2 explores only single solution path, successfully solving the problem by that path and without recalling very early steps.
5 Conclusion
We propose AutoMR, a framework that searches for query-aware meta-reasoning skeleton to guide LLM reasoning. By formulating meta-reasoning as a search problem over DAG-based search space, AutoMR covers skeletons in prior works and can capture intricate logical dependencies among reasoning steps. AutoMR designs a dynamic skeleton sampling algorithm that can derive any skeleton in search space within minimal additional computation overhead, and make skeleton adaptable to evolving base reasoning context, thus enabling efficient search. Experiments on math Q&A and general multiple-choice benchmark datasets demonstrate consistent improvements over existing meta reasoning methods.
Appendix A Implementation Details
A.1 Meta Reasoning Strategy Implementation
The functions of meta reasoning strategies is summarized in Table 3. We design maybe more than one prompts for each strategy and sample one randomly when sampling strategy for an edge. Some prompts are used only for certain tasks and we indicate them in parentheses after the prompt. The prompts of all meta level strategies are as follows.
Table 3: Meta reasoning strategies.
Strategy Function
Next Reason to next step.
Reflect Reflect previous reasoning steps
Explore Inspire divergent thinking
Decompose Decompose current query and propose sub-question.
Summarize Summarize previous reasoning steps.
Recall Recall related knowledge or previous steps about problem.
Answer Give answer and end current reasoning path.
Prompt for Next Next, Then, Now, let me move on to the next step.
Prompt for Reflect Let me consider what part of the reasoning feels least certain, and how can it be examined. Wait, let me think if there anything missing in the current reasoning. Let me think does the current line of thought have any error.
Prompt for Explore Let me consider which direction of thinking I should explore. Let me think what potential strategy has not yet been considered that could be the next solution path. Let me think what possible solution could be tried next.
Prompt for Decompose This question is a bit complex, let me think how to decompose it into sub-questions that I can solve. The question feels too broad, let me think what smaller version could I tackle first. Let me think if I can express the problem in terms of simpler components or modules. Let me consider the options one by one. (General multi-choice)
Prompt for Summarize Let me summarize what have I established so far. Let me summarize the current state of reasoning process, what’s known, unknown, and assumed? Let me consider if I can captures the essence of the reasoning so far with single sentence.
Prompt for Recall Let me think if I have encountered similar problems or if learned knowledge and previous intermediate step can be used here. Let me think what prior reasoning steps are directly relevant here or this question connect to earlier results. (Math Q&A). Let me recall which theorems, rules, or principles from earlier knowledge is related to this question. (General multi-choice).
Prompt for Answer Let me give the answer according to current reasoning context.
A.2 Meta Reasoning Strategy Sampling
We implement an MLP model to sample strategy for edge $(j,i)$ from $n_{j}$ to $n_{i}$ by taking representations of potential predecessor node content $c_{j}$, already sampled strategy $s_{>j,i}$ and current base reasoning context composed of all node content in partial skeleton $c_{:i-1}$.
Specifically, we maintain a learnable embedding layer to map each strategy $s\in{\mathcal{S}}\cup\{\textit{zero}\}$ to a dense embedding. For each node content $c$, we save the mean of “last hidden state” of the $c$ as semantic representation of the node content. “Last hidden state” is byproduct of LLM inference process for token distribution when generating each token, requiring no extra LLM invocation.
Finally, we build input for MLP according to $\text{Concat}([e(c_{j}),\text{Mean}(e(s_{>j,i})),\text{Mean}(e(c_{:i-1}))])$, where $\text{Concat}(\cdot)$ means concatenate vectors and $\text{Mean}(\cdot)$ means calculate the mean of vectors. We use $\text{Softmax}(\cdot)$ to process output of MLP and give the distribution of $s_{(j,i)}\in{\mathcal{S}}\cup\{\textit{zero}\}$.
A.3 Overall Search Algorithm
We show the overall search algorithm in Algorithm 2. We set implement $N$ as 8, $M$ as 16 and learning rate $\eta$ to $5\times 10^{-4}$ during search for both tasks. We refer to previous works [zhuge2024gptswarm, zhang2025multiagent, logicrl, reinforce++, cheng2020instanas], implement techniques such as gradient clipping, to improve the stability and convergence rate of search algorithm. See our code for implementation details. We implement a rule-based $r$ by exactly matching final answer $\hat{a}=\text{LLM}(q,\alpha_{q})$ given by LLM with ground-truth $a$ from dataset. Specifically,
$r(a,\text{LLM}(q,\alpha_{q}))=\begin{cases}1,&\text{if }\text{LLM}(q,\alpha_{q})=a,\\ -1,&\text{if }\text{LLM}(q,\alpha_{q})\neq a.\end{cases}$
Algorithm 2 Overall Search Algorithm
0: Dataset $\mathcal{D}$}, learning rate $\eta$
0: Trained $\theta$
1: Initialize $\theta$ randomly
2:while not convergence do
3:  Sample a batch $\{q_{1},q_{2},...,q_{N}\}$ from $\mathcal{D}$}
4:  Sample $\{\alpha_{q_{i}}^{1},\alpha_{q_{i}}^{2},...,\alpha_{q_{i}}^{M}\}$ for each $q_{i}$ with Algorithm 1
5:  $\theta\leftarrow\theta+\frac{\eta}{MN}\sum\limits_{i=1}^{N}\sum\limits_{j=1}^{M}[r(a_{i},\text{LLM}(q_{i},\alpha_{q_{i}}^{j}))\nabla_{\theta}\log P_{\theta}(\alpha_{q_{i}}^{j}|q_{i})]$
6:end while
7:return $\theta$
Appendix B Theoretical Analysis
B.1 Proof of Proposition 1
Proof.
We prove each case by construction.
Sequential. A sequential structure is defined as an ordered set of noes ${\mathcal{V}}=\{v_{1},\dots,v_{k}\}$ with edges
${\mathcal{E}}=\{(i,i+1)\mid 1\leq i\leq k-1\},$
and $\tau((i,i+1)\in{\mathcal{S}}$ for each $i$. Clearly, $v_{1}$ is the unique source ($\deg^{-}(v_{1})=0$ and $\deg^{-}(v)=1$ for all $v\neq v_{1}$), and $G$ is acyclic since edges only connect $v_{i}\to v_{i+1}$. Hence $({\mathcal{V}},{\mathcal{E}},{\mathcal{S}},\tau)$ is a single-source edge-heterogeneous DAG.
Tree. A tree is a rooted directed graph $G=({\mathcal{V}},{\mathcal{E}},{\mathcal{S}},\tau)$ such that:
$\exists!\ r\in{\mathcal{V}}\ \text{with }\deg^{-}(r)=0,\quad\forall v\in{\mathcal{V}}\setminus\{r\},\ \deg^{-}(v)=1.$
By definition, a rooted tree has no directed cycles and admits a unique source $r$. Since $\tau:{\mathcal{E}}\to{\mathcal{S}}$ can assign arbitrary heterogeneous edge types, $({\mathcal{V}},{\mathcal{E}},{\mathcal{S}},\tau)$ is a single-source edge-heterogeneous DAG.
Parallel. A parallel structure is defined by a common entry node $s$ and a family of disjoint branches
$\mathcal{B}=\{B_{1},\dots,B_{m}\},\quad B_{i}=({\mathcal{V}}_{i},{\mathcal{E}}_{i},{\mathcal{S}},\tau|_{{\mathcal{E}}_{i}}),$
where $s\in{\mathcal{V}}$ and for each $i$ we have $(s,u)\in{\mathcal{E}}$ with $u\in{\mathcal{V}}_{i}$ the root of branch $B_{i}$. Thus the overall structure is
${\mathcal{V}}=\{s\}\cup\bigcup_{i=1}^{m}{\mathcal{V}}_{i},\quad{\mathcal{E}}=\bigcup_{i=1}^{m}\big(\{(s,u_{i})\}\cup{\mathcal{E}}_{i}\big).$
This is precisely a rooted tree with root $s$ and subtrees $B_{i}$ attached as children. Therefore, a parallel structure is a special case of a tree, and hence also a single-source edge-heterogeneous DAG.
Since sequential, tree, and parallel (as a special case of tree) all admit representations $({\mathcal{V}},{\mathcal{E}},{\mathcal{S}},\tau)$ that satisfy (i) unique source, (ii) acyclicity, and (iii) heterogeneous edge labels, they are all contained in the class of single-source edge-heterogeneous DAGs. ∎
B.2 Proof of Proposition 2
We first prove that Algorithm 1 can cover any skeleton $\alpha\in{\mathcal{A}}$ and then analyze the time complexity.
Proof.
Since $\alpha$ is acyclic, by a standard result there exists a topological ordering of its vertices. That is, there exists a permutation $\pi=(n_{1},n_{2},\dots,n_{|{\mathcal{V}}|})$ of $\mathcal{V}$} such that for every edge $(u\to w)\in{\mathcal{E}}$ we have $u$ appears earlier than $w$ in $\pi$.
Use this topological order $\pi$ as the insertion order in the append-only construction: add nodes in order $n_{1},n_{2},\ldots,n_{|}{\mathcal{V}}|$. When adding $n_{t}$, consider all previously added nodes $\{n_{1},\dots,n_{t-1}\}$. Because $\pi$ is a topological order, every edge in $\mathcal{E}$} that is incident to $n_{t}$ from earlier nodes is of the form $n_{i}\to n_{t}$ with $i<t$; there are no edges from $n_{t}$ back to any already-added node. Therefore, by choosing exactly those forward edges $\{(n_{i}\to v_{t})\in{\mathcal{E}}\mid i<t\}$ at step $t$, we add precisely the edges of $\alpha$ that end at $n_{t}$.
Applying this procedure for $t=1,\dots,n$ adds all and only the edges of $\alpha$. Hence the append-only construction, with insertion order equal to any topological order of $\alpha$ and with edge choices equal to the edges of $\alpha$, reproduces $\alpha$ exactly. ∎
Besides invoking the LLM to generate textual reasoning content, Algorithm 1 requires at most $O(|{\mathcal{V}}|^{2})$ sampling process for reasoning steps count $|{\mathcal{V}}|$ with two layers of “for” loop, where each sampling process corresponds to a single MLP call.
Let $\mathcal{B}$} denote token budget of the generated reasoning content. Since Algorithm 1 introduces no additional LLM calls as analyzed in Section 3.2, the time complexity of LLM invocation remains $O({\mathcal{B}}^{2})$.
In practice, the reasoning step count $|{\mathcal{V}}|$ is roughly proportional to $\mathcal{B}$}, but typically $|{\mathcal{V}}|\ll{\mathcal{B}}$, as each reasoning step consists of many tokens.
Furthermore, the computational cost of MLP inference is negligible compared with the layered blocks of the LLM. Therefore, AutoMR introduces only minimal additional computational overhead relative to naive LLM reasoning.
Appendix C Experiment Details
C.1 Baseline Implementation
The system prompt and answer extraction code for math Q&A problem is referred to a open-source repository openr  [^1] https://github.com/openreasoner/openr. The system prompt and answer extraction code for general multiple-choice problem is referred to the original MMLU-Pro repository  [^2] https://github.com/TIGER-AI-Lab/MMLU-Pro.
For all baselines, we implement with Qwen and LLaMA as base model rather than the LLM used in their original paper for fair comparison.
  • MRP. MRP does not have open-source code, but provides prompt in original paper. We follow the paper to implement MRP.
  • Meta-Reasoner. Meta-Reasoner does not have open-source code, but provides prompt, pseudo code and detailed description in original paper. We follow the paper to implement Meta-Reasoner.
  • rStar. We implement rStar with it open-source code  [^3] https://github.com/zhentingqi/rStar.
  • MaAS. We implement MaAS with it open-source code  [^4] https://github.com/bingreeky/MaAS.
  • RS. Referring to previous works [randomsearch, darts], we sample 48 architectures from search space randomly. Then we validate these architectures on training set to select the one with highest accuracy. With the selected architecture, we report its accuracy on test set.
  • QI. Referring to previous works [darts, zhuge2024gptswarm], we do not use an MLP which takes reasoning context as input and output meta strategy distribution, but model the strategy distribution of each edge in search space without condition. We optimize the distribution with the same estimation of policy gradient with REINFORCE as in Equation 4. For all queries in test set, we sample only one skeleton to process all of them.
  • CA. Referring to previous works [cheng2020instanas, zhang2025multiagent], we use an MLP which takes semantic embedding of queries and meta reasoning strategies existing in skeleton as input to sample strategy for edges, rather than based on base reasoning context. For each query in test set, we sample a complete skeleton before inference and then reason for the query guided by the complete skeleton.
C.2 Datasets Details
For training set, we use MATH  [^5] https://github.com/hendrycks/math training split composed of 5053 query-answer pairs and MMLU-Pro  [^6] https://github.com/TIGER-AI-Lab/MMLU-Pro training split composed of 70 query-answer pairs. For testing set, we use GSM8K  [^7] https://github.com/openai/grade-school-math, MATH-500  [^8] https://huggingface.co/datasets/HuggingFaceH4/MATH-500, AMC  [^9] https://huggingface.co/datasets/AI-MO/aimo-validation-amc, Olympiad  [^10] https://github.com/OpenBMB/OlympiadBench and four subset (Science, Humanities, Social and Other) of MMLU-Pro. We summarize the statistics of dataset in Tabel 4.
Table 4: Dataset Statistics.
Domain # Train Dataset # Test Description
Math Q&A 5053 GSM8K 1319 Grade school math.
MATH-500 500 High school math.
AMC 83 High school competition math.
Olympiad 674 Olympiad-level math competition.
General Multi-Choice 70 Science 5345 Physic, chemistry, biology, etc.
Humanities 1981 Philosophy, history and law.
Social 2431 psychology, business and economics.
Other 924 Other topics
C.3 Additional Results on Larger LLM
Besides Qwen2.5-3B-Inst and LLaMA3.2-3B-Inst used in Section 4.1 with token budget 1024, we conduct additional experiments to evaluate AutoMR based on larger long-cot LLM, including Qwen3-8B and Qwen3-14B (both enabling thinking mode, refer to Qwen3 homepage  [^11] https://qwenlm.github.io/blog/qwen3/ for more details about thinking mode) with larger token budget 8192. Additionally, we also evaluate on Pangu-7B with token budget 4096 on Ascend. We evaluate AutoMR and two meta reasoning methods that performs well in Table 1 on two datasets (MATH500 and Olympiad).
The results are showed in Table 5, where Base means the thinking mode of Qwen3 and Pangu7B with CoT. These results show that AutoMR still achieves performance gains when scaling to larger LLMs with larger token budget, compared with naive LLM thinking and previous meta reasoning methods, demonstrating the scaling potential of AutoMR.
Table 5: Additional results on Qwen3-8B, Qwen3-14B and Pangu-7B.
Method MATH-500 Olympiad
Pangu-7B Qwen3-8B Qwen3-14B Pangu-7B Qwen3-8B Qwen3-14B
Base (S) 75.4 80.8 84.2 48.5 47.0 49.0
Meta-Reasoner(S) - 81.6 85.6 - 49.9 50.4
rStar (T) - 83.4 84.8 - 48.9 50.9
AutoMR (G) 77.4↑2.0 86.6↑5.8 87.2↑3.0 49.2↑0.7 53.9↑6.9 54.5↑5.5